OpenStudy (anonymous):

The variables x and y vary inversely. Which equation relates x and y when x = -1 and y = -9?

OpenStudy (ganpat):

\[x \alpha (1/y) = x = K/ y\] This can be equation.. SO now put x & y values in equation.. -1 = K/ -9 K = 9 So the equation would be , x = 9/ y...

OpenStudy (anonymous):

Find the sum of the series: \[\sum_{i=1}^{4}6i-5\]